### v1

# Adaptive Super-Resolution Imaging Without Prior Knowledge Using a Programmable Spatial-Mode Sorter

Abstract
We consider an imaging system tasked with estimating the angular distance between two incoherently-emitting sub-Rayleigh-separated point sources, without any prior knowledge of the centroid of the constellation and with a fixed collected-photon budget. It was shown theoretically that splitting the optical recording time into two stages—focal-plane direct imaging to obtain a pre-estimate of the centroid, and using that estimate to center a spatial-mode sorter followed by photon detection of the sorted modes—can achieve 10 to 100 times lower mean squared error in estimating the separation [1]. In this paper, we demonstrate this in proof-of-concept, using a programmable mode sorter we have built using multi-plane light conversion (MPLC) using a reflective spatial-light modulator (SLM) in an emulated experiment where we use a single coherent source to characterize the MPLC to electronically piece together the signature from two closely-separated quasi-monochromatic incoherent emitters.
journal: opticajournal
1 Introduction
Adaptive optical imaging is an increasingly prevalent technology, enabling automated imaging systems in diverse fields such as biomedicine [2], astronomy [3], remote sensing, and autonomous navigation. Adaptive imaging refers to the manipulation of optical hardware during an acquisition, often in a multi-stage procedure, such that processed data from initial stage(s) determine the imaging configuration used in subsequent stage(s). While dynamic scenes often necessitate adaptive imaging schemes, optical adaptation is also a powerful tool for static imaging, allowing for preliminary acquisition of "nuisance" information that is not critical to the final imaging result but helps inform the appropriate manipulations that will best perform the desired imaging task [4, 5, 1]
For both adaptive and non-adaptive imaging, directly passing light from the scene through conventional optics like lenses and mirrors before reaching a focal-plane detector array constitutes the most widespread method. While effective for many purposes, this "direct imaging" approach encounters a resolution constraint often referred to as the Rayleigh limit ($\theta_{\rm{res}}$) [6], an angular distance scale beyond which the features of the scene become exceedingly hard to resolve. The limitation in resolution stems from diffraction effects arising due to the finite aperture of the imaging system, as illustrated in Fig.1. For a telescope-like system with a circular aperture, this limit is defined as follows:
$\theta_{\rm{res}}=1.22\times\frac{\lambda\times f}{D},$ (1)
where $\lambda$ is the wavelength of the incoming light, $f$ is the focal length of the optical system, and $D$ is the diameter of the entrance pupil of the optical system, which is dictated by the finite circular aperture. An analogous diffraction limit is encountered in microscopy due to the finite numerical aperture of a circular objective lens.
Figure 1: Example image plane intensity distribution from two point sources with separation $\theta$, centroid $\phi$, aperture diameter $D$, focal length $f$, and PSF width $\sigma$.
While sub-diffraction imaging can be done in the near field [7], with controllable fluorescent probes [8, 9, 10, 11], and via structrured optical excitation [12], fully passive optical super-resolution was made possible by a recent class of imaging receivers known as spatial mode demultiplexing (SPADE) [13]. SPADE uses physical linear optical processes such as multi-path interferometry [14] or holography [15] to couple specific orthogonal spatial modes of the incoming optical field to corresponding intensity detectors. References present various experimental demonstrations showcasing resolution enhancement below the Rayleigh limit in the transverse distance estimation problem through the utilization of spatial mode demultiplexing to classify Hermite-Gauss (HG) modes, prior to detection. Recently, multi-plane light conversion (MPLC), which uses a succession of phase masks to transform the optical wavefront [16], has emerged as the most promising SPADE implementation. Notably, a commercially available MPLC device that demultiplexes a fixed set of 2D spatial modes (Hermite Gauss modes) has spurred experimental progress for several recently demonstrated sub-diffraction imaging tasks [17, 18, 19, 20].
Neither SPADE nor MPLC are constrained to a particular set of spatial modes, and a device that can update operational spatial mode basis would be a perfect fit for adaptive super-resolution imaging. Several theoretical works have proposed using mid-acquisition adaptation to extend passive super-resolution techniques to more complex imaging tasks [1, 21, 22, 23]. Experimental progress in adaptive mode sorting is limited to date, with some systems demonstrating limited adaptability [24]. A fully adaptive MPLC implementation could perform highly complex imaging tasks in the presence of nuisance parameters.
Our experiment attempts to complete a more complicated task, demonstrating resolution enhancement below the Rayleigh limit in the transverse distance estimation problem without prior knowledge of the locations of the two incoherent sources or the centroid of the scene [1]. In addition, we employ a reconfigurable mode sorter, potentially applicable in more challenging localization tasks in subsequent iterations [23].
2 Main Results
Our experiment primarily aims to estimate the separation ($\theta$) between two identical incoherent point sources with an unknown centroid, illustrated in Figure 1. To accomplish this, we adopted the strategy that was first hinted in [13] and fully outlined in [1], which suggests dividing photon resources between direct detection for centroid estimation and binary mode sorting for separation estimation. Direct detection has the advantage of shift invariance, allowing it to identify the location of the source pair over a wide field of view, and also is close to an optimal measurement of the object centroid [25]. Our second-stage SPADE measurement has been shown to achieve the resolution quantum limit for estimating small separations between point sources when the mode sorter is nearly perfectly aligned [13].
The principal findings, expounded upon in the subsequent results section, reveal notable enhancements. Specifically, our hybrid approach demonstrates a reduction in variance of separation estimation for separations of up to 0.3 Rayleigh units, as well as a decrease in mean squared error of separation estimation for separations of up to 0.2 Rayleigh units when contrasted with direct imaging. The estimation of separations is executed using maximum likelihood estimators applied to both the direct imaging data and the sorted data.
The findings outlined in this paper serve as a proof of concept. It is important to note that the experimental results are partly emulated, given that we did not employ two distinct incoherent sources simultaneously. Instead, data acquisition was conducted sequentially from individual sources and subsequently merged computationally. Additionally, we took compensatory measures to account for losses attributable to our mode sorter, as detailed in the supplementary materials. This ensured parity in photon count between the two compared methods: direct imaging and the hybrid approach of half direct imaging and half mode sorting.
3 System Overview
Figure 2: (a) sample phase masks that would be displayed on the SLM, where SLM 1 acts as the object under test and the following three masks act as the MPLC masks. (b) Experimental setup.
We constructed a system that is able to acquire direct imaging data from the scene under test, estimate the scene’s centroid, align itself to the found centroid, and sort the appropriate orthogonal mode basis.
It is important to note that this setup is capable of more than separation estimation between 2 point sources. However, in this paper, we will focus on this specific application.
3.1 System walk through
Our system consists of 4, main, subsystems:
Illumination: Our 532 nm laser source (Edmund Optics 10 mW DPSS) first goes through a linear polarizer in order to align its polarization to the polarization required by the SLM. A spatial filter is then used to eliminate the higher spatial frequencies of the beam and produce a clean, uniform, Gaussian beam. The filtered beam is then collimated and directed to the SLM.
SLM: In our system we use a liquid crystal based SLM (Holoeye PLUTO-2.1-VIS-016), which will cause aberrations in the system if not accounted for (See supplementary material). As can be seen in Fig.2(a), four square regions are displayed on the SLM. The first region is used to encode a phase-modulated object, and the other three are used as phase masks created by our algorithm, which was based on [16], forming the MPLC-based optical transformation that will ideally direct each orthogonal input mode to a respective, pre-set, location on the detector. On each square, we applied a blazed grating and used the first diffraction order rather than the zeroth order. The zeroth order is noisier due to the reflection of unmodulated light, while the first diffraction order is lossier. We use the latter as our focus is to minimize the crosstalk in mode sorting. Note that due to the SLM’s programmability and the fact that our MPLC- based code works for every orthogonal mode set, we achieve a reconfigurable mode sorter.
4f system: As seen in Fig.2(b), the light goes through a 4f system right after the object is encoded at the first bounce from the SLM. The reason that we need the 4f system is to blur our phase-encoded object. By placing a pinhole (PH2 in Fig.2(b)) at the Fourier plane of the 4f system we effectively create a low pass filter, and by changing the size of PH2 (in our experiment we used a 200 microns pinhole, we can change the width of the PSF. This allows us to experiment with super-resolution imaging even though our object itself is not small enough. Due to the nature of the 4f system, the conjugate planes at the start and the end of the 4f system will be identical (if we take PH2 out) and therefore the phase-encoded object is going to be presented to the remaining SLM planes for processing.
Detection: As shown in Fig.2(b), we use a beam splitter (BS) to be able to access both a direct image and the mode demultiplexed detection data. Note that only one of the outputs is going to reach the camera at a time and this will be decided by the flip mirror (M8 in Fig.2(b)). Each of these outputs is detected using the photon counting feature of an electron-multiplying charge-coupled device (EMCCD) camera (Andor iXon 897).
A more detailed system walk-through can be found in [26].
3.2 Zernike Modes Sorting Characterization
The ideal modes to sort for our experiment’s purposes are the Zernike (or airy disc) mode family, since it is the PSF-adapted mode for hard circular aperture [27]. To ensure the feasibility of the experiment we tested how well our mode sorter sorts the Zernike modes of interest. To conduct that test we displayed a point source as the object (SLM 1 in Fig.2a) and swept it across the whole scene. As can be seen in Fig.3, the experimental data from the point source sweep closely matches the theoretical results, especially when the deviation off-axis is small. When the object is exactly on axis (where theoretically we expect all the photons to be in the Zernike00 mode) we saw only 0.89% cross-talk (definition of cross-talk further discussed in the supplementary material).
Note that while in [1] binary mode sorting is discussed, in Fig.3 we can see sorting of 3 modes. In our experiment, we still tried to imitate the binary mode sorting case. However, as the separation between the point sources increased we saw more contribution from higher order mode (as expected). To mitigate this contribution and ensure that it does not corrupt our data we decided to sort 3 modes (Zernike00, Zernike01, and Zernike02). The objective of sorting the third mode (Zernike02) was to ensure that the photons that live in the higher-order modes (especially when the separation increases) are directed there rather than to the other two spots which would corrupt the data. The other two spots (Zernike00 and Zernike01) were then used to evaluate the binary mode sorter’s performance in super-resolution imaging.
Figure 3: Calibration plot of theoretical (solid lines) and experimental (circles) intensities found in each mode with respect to the distance of the point source away from the known center. Amplitudes of the three orthogonal spatial modes that are sorted by the MPLC are shown on the left. Below the plot are the raw experimental images with the regions designated for each mode shown in circles, where each image corresponds to the indicated distance away from the system’s axis. Note: we report crosstalk value of 0.9831 between the fundamental mode and the higher order modes when the point source transition=0.
4 The Experiment
4.1 Experiment Overview
Our experiment is divided into 2 different stages, and its objective is to prove the capability of our mode sorting system to achieve super-resolution capabilities with relaxed prior information. More specifically, we wanted to exceed the Rayleigh limit [6] when estimating the separation between two point sources without knowing their centroid. The way in which we conducted our 2-stage experiment can be seen in [1]. To summarize, first, direct imaging data was acquired, from this data the centroid of the two point sources can be approximated (Section 4A).
Once the centroid is approximated, our MPLC-based algorithm creates the appropriate phase masks for the found centroid to perform the mode sorting. This generation of special phase masks essentially aligns our system to the scene and eliminates any need for physical manipulations to the system. Finally, using these specially created phase masks, mode sorting is performed. From this data (sorted data), we estimate the separation between the two point sources using maximum likelihood estimation as part of the data processing. In this experiment, we tested 12 different separations (from 0.1 Rayleigh units of separation to 1.2 Rayleigh units of separation in increments of 0.1 Rayleigh units) and had 200 data points for each separation. Each data point had a randomly selected centroid within the field of view of our system.
All the data in this experiment was taken using the photon counting mode in our EMCCD and the total number of photon arrivals at each data point (which is a combination of two images) was aimed to be 10,000. The laser power going into the system was calibrated accordingly while ensuring a low likelihood of multiple photons events. (See supplementary materials section D). Note: We conducted the experiment using a laser (coherent source), while for the experiment to correctly correlate the theory of super-resolution for two incoherent point sources, the light from the two point sources cannot interfere with each other. We therefore acquired data from only one point source at a time and then combined the two images computationally.
4.2 First Stage
Figure 4: Example output from direct imaging with different separations. As the separation increases it is easier to notice that two point sources are present.
In the first stage of this experiment, we acquired direct imaging data by sweeping the point source (20x20 pixels on the SLM) across the full field of view (200 pixels vertically) 5 times. The reason that we did it 5 times is to be able to combine images from different sweeps which will supposedly make it more difficult to estimate both the centroid and the separation and will make the scenario more realistic. Once all the data was acquired, we randomly chose 200 different centroids and for each random centroid we chose the appropriate frame from a random data set (each one of the 5 point-source sweeps is considered one data set). After the chosen frames were combined into one image (an example of that with all tested separations is shown in Fig.4), we estimated the centroid using the center of mass method.
After the centroid was estimated, it was fed into to the MPLC based, phase mask generation algorithm which created the appropriate phase masks to sort the wanted modes where the estimated centroid is set to be the center of the scene. A total of 2400 (12 separations and 200 data points for each separation) centroids and phase mask sets were generated.
4.3 Second Stage
Figure 5: Example output from sorting. The left spot represents the relative presence of the 0th order mode, the right spot represents the relative presence of the 1st order mode, and the upper left spot (dim and only present in higher separations) represents the presence of the 2nd order mode.
The second stage of the experiment focused on acquiring the mode sorting data (shown in Fig.5). In this stage, we imaged the point sources (with the mode sorter) at a specified location, which is the same randomly chosen location discussed as part of the first stage of the experiment using the appropriate phase masks set that was generated using the centroid estimation from the matching direct imaging data. This was done 4800 times (each data point required 2 point sources, and we did it 200 times for the 12 separations that were tested).
5 Data Processing
5.1 Centroid Estimation
To estimate the centroid we used the direct imaging data (as shown in Fig.4) from different trials, as mentioned in the First Stage sub-section. Once we had the computationally combined image we used the center of mass method to estimate the centroid.
Our detector had 512x512 pixels so the center of the camera was set to be the zero position. Each pixel to the left for the x centroid ($\phi_{x}$), or down for the y centroid ($\phi_{y}$) was assigned a value from -256-0 accordingly. Similarly, each pixel to the right for $\phi_{x}$ or up for $\phi_{y}$ was assigned a value from 0-256 accordingly. This gave us two vectors:
$V_{x},V_{y}=[-255:256]\times PixelPitch$. Using these vectors in a mesh grid ($V_{x}\rightarrow M_{x_{i,j}}$, $V_{y}\rightarrow M_{y_{i,j}}$) along with the acquired data ($Im_{i,j}$) we can calculate the estimate of the x and y centroids as such:
$\phi_{x}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{x_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}},\rm{and}$ (2a)
$\phi_{y}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{y_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}}.$ (2b)
This method is valid in our case since we used equally intense point sources, however, this approach will need to be modified if point sources of different intensities are under test.
5.2 Separation Estimation- Direct Imaging
Since in our experiment we used a finite circular hard aperture we can describe the intensity pattern resulting from a point source as an airy disk:
$S(x,y)=\left(\frac{2\times J\left(\frac{\pi\times\sqrt{x^{2}+y^{2}}\times D_{\rm{ph}}}{\lambda\times f}\right)}{\frac{\pi\times\sqrt{x^{2}+y^{2}}\times D_{\rm{ph}}}{\lambda\times f}}\right)^{2}\rm{,}$ (3)
where $J(.)$ is the Bessel function, $D_{\rm{ph}}$ is the diameter of the circular aperture, $f$ is the focal length of the lens used in the system, and $\lambda$ is the wavelength.
Because in this experiment we have two point sources separated in 1D we can describe the intensity pattern on the detector as such:
$\Psi(x,y)=S(x+\phi_{x}+\frac{\theta}{2},y+\phi_{y})+S(x+\phi_{x}-\frac{\theta}{2},y+\phi_{y}).$ (4)
Using the intensity pattern on the detector as well as the estimated centroid we can write a likelihood function:
$P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)=\prod_{i=1}^{n_{1}}\Psi(x_{i},y),$ (5)
where $\theta$ is the separation between the point source, $x_{i}$ stands for the position of the $i^{\text{th}}$ detected photon, and $n_{1}$ stands for the total number of photons detected in the direct imaging stage.
Now, by maximizing the likelihood function we get our estimation for the point-sources separation :
$\hat{\theta}_{\rm{direct}}=\mathrm{argmax}_{\theta}(P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)).$ (6)
5.3 Separation Estimation- Sorting
In this paper, we present two ways in which we processed our sorting data: 1) Maximum likelihood- This is the way that we used to characterize the theoretical performance of the system as this method should give us the most accurate estimation of the parameter of interest, $\theta$ when the distribution of the data follows the model exactly. 2) Comparison to characterization- Since our system may have non-idealities, such as slight misalignments, there is a chance that the distribution will not exactly follow the theoretical model. Therefore, in this method, we compared our raw data to a performance characterization of our system, as seen in Fig. 3.
5.3.1 Maximum Likelihood
In our experiment, we sorted Zernike or Airy disk modes. However, for simplicity, and due to the similarity of their plots in the super-resolution regime we matched our binary SPADE photon counting results to the probability of a photon landing in the 0th and 1st Hermite-Gauss modes as a function of the separation, $\theta$, which can be described as such:
$\psi_{0}=\frac{\exp{\frac{-\theta^{2}}{\sigma^{2}}}}{\max(\exp{\frac{-\theta^{2}}{\sigma^{2}}})},\rm{and}$ (7a)
$\psi_{1}\approx 1-\psi_{0},$ (7b)
where:
$\sigma=1.22\times\frac{\lambda\times f}{D_{\rm{ph}}}.$ (8)
$\psi_{1}$ can be described as it is in eq.7b because we are doing binary SPADE and therefore, all the photons that will not be counted in the $\psi_{0}$ spot, will be counted from the $\psi_{1}$ spot.
It is important to note that when applying this model to our experimental cross-talk ($C$) (see supplementary material section A for the cross-talk definition) was taken into account such that $\psi_{0}$ and $\psi_{1}$ become:
$\psi_{0}=C\times\frac{\exp{\frac{-\theta^{2}}{\sigma^{2}}}}{\max(\exp{\frac{-\theta^{2}}{\sigma^{2}}})},\rm{and}$ (9a)
$\psi_{1}=\exp{\frac{-\theta^{2}}{\sigma^{2}}}\times 2\times(\frac{\theta}{\sigma})^{2}.$ (9b)
Using the description of photon counts expectation as a function of $\theta$ we can describe the sorting likelihood function $P_{B}$ as a binomial distribution:
$P_{B}(I_{0},n_{2}|\theta)=\begin{pmatrix}n_{2}\\ I_{0}\end{pmatrix}(\psi_{0})^{I_{0}}(\psi_{1})^{n_{2}-I_{0}},$ (10)
where $I_{0}$ is the number of photons detected in the $\psi_{0}$ spot, and $n_{2}$ stands for the total photons detected in the sorting stage.
Now, to estimate the most likely $\theta$ we maximize the likelihood function $P_{B}$:
$\hat{\theta}_{\rm{Sorting}}=\mathrm{argmax}_{\theta}(P_{B}(I_{0},n_{2}|\theta)).$ (11)
5.3.2 Comparison to Characterization
The exact number of photons that gets detected at every data point will vary. Because of that, rather than comparing the number of photons in each mode to our characterization we found the ratio between $\psi_{0}$ and $\psi_{1}$ from our data and matched it to the same ratio from our interpolated system’s characterization (Fig. 3). By matching the ratios we could then find what is the separation that is the most fitting for the data.
5.4 Separation Estimation- 2-Stage
As mentioned in section 3A, when we try to estimate $\theta$ using our mode sorter, we still use direct imaging for about half of the photons for the centroid estimation. In this method instead of just using these photons for centroid estimation we also use them for the separation estimation. This leads to a new likelihood function that utilizes all the photons collected:
$P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x},I_{0},n_{2}|\theta)=P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)\times P_{B}(I_{0},n_{2}|\theta).$ (12)
Therefore, the estimated $\theta$ using this method can be found by maximizing the above likelihood function:
$\hat{\theta}_{{\operatorname{2-stage}}}=\mathrm{argmax}_{\theta}(P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x},I_{0},n_{2}|\theta)).$ (13)
6 Results
The goal of this experiment was to prove the superiority of our adaptive mode sorter in separation estimation below the Rayleigh limit when compared to traditional direct imaging.
Figure 6: Monte-Carlo simulation that shows the mean squared error (MSE) of separation estimation using 10000 photons of direct imaging (blue circles), 5000 photons in direct imaging for centroid estimation, and 5000 photons for sorting (orange circles), and 5000 photons in direct imaging for centroid estimation and for separation estimation, and 5000 photons for sorting (yellow circles). Note that because in our simulation we assumed the absence of bias the plot for the MSE vs. the separation will look identical to the plot for variance vs. separation
Figure 7: a) Mean squared error of the separation estimation , b) Variance of the separation estimation using 10000 photons of direct imaging (blue circles), 5000 photons in direct imaging for centroid estimation and for separation estimation,5000 photons for sorting with the maximum likelihood method (orange circles), and 5000 photons for sorting with the comparison to characterization method (purple circles)
Before conducting the experiment, we ran a Monte Carlo simulation. The results of the simulation are promising as the separation estimation MSE in the deep sub-Rayleigh regime (below 0.5 Rayleigh units of separation) shows an improvement by a factor of up to 12.7. Note that in the Monte Carlo simulation, we assumed a perfect mode-sorter, with no crosstalk, (see supplementary information), a perfect detector (no noise). The simulation results can be seen in Fig. 6. In this figure, when looking at the simulation results when only using the sorting data for separation estimation (orange circles), one can notice a counter-intuitive phenomenon, the MSE increases with the separation. This cam be explained by the fact that at very small separations most of the Fisher information exists in the modes that we sort (Zernike00 and Zernike01) and as the separation increases there is more Fisher information that lives in higher order modes. Another way to think about it is by looking at fig.3. When the separation is very small, there are very few photons arriving at the Zernike01 spot and every change in that is easily detected. However, as the separation increases, small changes at the intensity of the Zernike01 spot become harder to detect and therefore making an error becomes more likely.
Fig.7a presents our experimental MSE results. One can see that in the deep sub-Rayleigh regime (up to 0.2 Rayleigh units of separation) our mode sorter is up to 2.56 times more accurate than experimental direct imaging and only 3.63 times less accurate than the Monte Carlo simulation which simulates a perfect detector.
In Fig.7b one can see that the variance of our mode sorter is lower than that of direct imaging up to 0.3 Rayleigh units of separation and is very close to it at 0.4 Rayleigh units of separation.
This tells us that there is an unexpected source of bias in our mode sorting data (this will be further discussed in the Supplementary Materials in section F). Some possible sources of error that may have degraded or biased our data are: crosstalk, although it was taken into account in the data processing, misalignment of the system, generally and with respect to the centroid, and shot noise in the detector. See supplementary information for further discussion of the sources of error.
7 Acknowledgments
This research was supported by the DARPA IAMBIC Program under Contract No. HR00112090128. The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.